OpenStudy (anonymous):

help for a medal and fan.... Which of the following is equivalent to (ab)4? a4b a4b4 ab4 4ab

OpenStudy (anonymous):

\[(a \times b)^4\] There's a power property that says \[(a \times b)^n = a^n \times b^n \] You just have to apply the property

OpenStudy (kamibug):

\(\Huge\bf\color{hotpink}{Welcome~to~Openstudy!~:)}\)

OpenStudy (anonymous):

i dont know what the property is @Umangiasd

OpenStudy (kamibug):

He said in his comment. ^^^ It's \[(a * b)^n = a^n * b^n\]

OpenStudy (anonymous):

its gonna be a4b4

OpenStudy (anonymous):

That's right